OpenStudy (anonymous):

The compound PtCl2(NH3)2 is effective as a treatment for some cancers. It is synthesized by the reaction shown in the equation K2PtCl4(aq) + 2 NH3(aq) → 2 KCl(aq) + PtCl2(NH3)2(aq). How much K2PtCl4 must react in order to produce 22.6 g of PtCl2(NH3)2? We have to answer it in units of mol. Help?

OpenStudy (aaronq):

So, you first need to convert the mass of the product to moles, then use those to trace back how much you need by using their stoichiometric ratio. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. --------------------------------------------------